# Bayesian Rank-Clustering

### v1

@@ Abstract
In a traditional analysis of ordinal comparison data, the goal is to infer an overall ranking of objects from best to worst with each object having a unique rank. However, the ranks of some objects may not be statistically distinguishable. This could happen due to insufficient data or to the true underlying abilities or qualities being equal for some objects. In such cases, practitioners may prefer an overall ranking where groups of objects are allowed to have equal ranks or to be rank-clustered. Existing models related to rank-clustering are limited by their inability to handle a variety of ordinal data types, to quantify uncertainty, or by the need to pre-specify the number and size of potential rank-clusters. We solve these limitations through the proposed Bayesian Rank-Clustered Bradley-Terry-Luce model. We allow for rank-clustering via parameter fusion by imposing a novel spike-and-slab prior on object-specific worth parameters in Bradley-Terry-Luce family of distributions for ordinal comparisons. We demonstrate the model on simulated and real datasets in survey analysis, elections, and sports.
@@ 1 Introduction
In a traditional analysis of ordinal data, we assume $I$ judges assess $J$ objects by providing ordinal preferences, $\Pi$. Each judge’s ordinal preferences, $\Pi_{i}$, may be provided in various forms, such as complete rankings, partial rankings, or pairwise comparisons among the available objects or some subset thereof. Standard statistical model families for ranking data (e.g., Mallows (Mallows, 1957), Bradley-Terry-Luce (Bradley and Terry, 1952; Plackett, 1975; Luce, 1959)) derive or estimate the rank of each object such that each object receives a unique rank. An overall ranking orders all objects from best to worst on the basis of their estimated rank. Analyses of this kind are used to rank candidates in an election given ranked choice votes (Gormley and Murphy, 2008; Mollica and Tardella, 2017), sports teams or players in a league using game outcomes (Tutz and Schauberger, 2015; Barrientos et al., 2023), or genes based on ordinal comparisons of genomic data (Eliseussen et al., 2023; Vitelli et al., 2018).
@@
However, requiring estimated ranks to be unique is not always useful or appropriate. Specifically, some objects may be equal or indistinguishable in quality or ability at the population level. In such cases, we refer to the objects as being rank-clustered. For example, consider one situation of an election in which 2 candidates, both of the same political party, are running for an office. If voters express their preferences solely on the basis of party, the candidates are inherently equal in quality; the rational voter would randomly select a candidate to vote for. In another case, when the number of votes cast is small, estimated ranks assigned to each candidate could exhibit substantial uncertainty, suggesting the candidates are indistinguishable in quality based on the limited number of observed votes. In such situations, allowing for rank-clusters may improve interpretability and aid accurate inference, prediction, and decision-making when analyzing ordinal preferences.
@@
In this paper, we propose a Bayesian framework for ordinal data analysis that estimates an overall ranking of objects with rank-clusters. Specifically, we choose to model observed rank via the Bradley-Terry-Luce (BTL) family of distributions which permits analysis of ordinal preferences in many forms, such as complete rankings, partial rankings, pairwise comparisons, and groupwise comparisons. To induce rank clusters, we place a novel spike-and-slab prior that partitions continuous parameters, proposed herein, on the object-specific parameters of BTL distributions. In contrast to existing work related to rank-clustering in the literature, our model does not require the parameter order nor the number or size of rank-clusters to be known in advance. Instead, these quantities are treated as random variables and estimated simultaneously, and their corresponding uncertainty is included in the resulting inferences. In this paper, we develop a computationally-efficient Gibbs sampler for estimation and apply the model to real and simulated data.
@@
The rest of this work is organized as follows. We first review relevant literature related to rank-clustering in Section 2. Then, we propose the Partition-based Spike-and-Slab Fusion prior and apply it to a BTL model for ordinal data in Section 3. We develop a computationally-efficient Gibbs sampler based on reversible jump Markov chain Monte Carlo in Section 4 and demonstrate its accuracy on simulated data. To demonstrate a wide variety of methodological values of our proposed framework, Section 5 applies the model to four real datasets: (i) complete rankings of sushi preferences provided by Japanese adults, (ii) partial rankings of 2021 Minneapolis mayoral candidates expressed by voters in a ranked choice election, (iii) complete and partial rankings of policy options from Eurobarometer 34.1, a survey which measures various European attitudes, and (iv) pairwise basketball game outcomes from the 2020-21 season of the National Basketball Association. We conclude with a brief discussion in Section 6.
@@ 2 Background
Before considering existing literature related to this paper, it is helpful to review some basic terminology and notation for ranking data. Rankings are a type of ordinal preference data that denote a relative ordering of objects from best to worst (potentially allowing ties). We use the operator $\prec$ to denote a strict ordering of two objects; e.g., $A\prec B$ states that object $A$ is strictly preferred to $B$. In a ranking, an object’s rank is the place it receives in the ranking. [^1] Although some authors have drawn a distinction between the terms “ranking” and “ordering,” in this paper we choose to use solely the former in accordance with its popular usage. Rankings arise in different forms. Given a collection of objects, a ranking is called complete when all objects are ranked. In contrast, a ranking is called partial when all objects were considered, but only a subset of the most-preferred are ranked (e.g., a top-$5$ ranking). In a partial ranking, we assume that unranked objects are less-preferred than those ranked, but also that the preference order among the unranked objects is unknown. Next, we call a ranking incomplete when a judge is asked only to rank a subset of the complete collection of objects. In incomplete rankings, no information can be gleaned regarding objects not considered. For example, if a voter is asked by an election pollster to rank candidates from a single political party, the ranking should provide no information regarding their preferences on candidates from other parties. We call incomplete rankings involving two objects (candidates in the above example) a pairwise comparison, and incomplete rankings involving more than two objects a groupwise comparison. Rankings may be both partial and incomplete; e.g., a top-3 ranking of mayoral candidates only among the 5 candidates from a specific political party.
In the rest of this section, we will briefly review methods for estimating rank-clusters based on the Bradley-Terry-Luce and Mallows families of ordinal data models in turn. For a more thorough review of these standard model families, see e.g., Marden (1996); Alvo and Yu (2014).
@@ 2.1 Methods based on Bradley-Terry-Luce Distributions
Since BTL distributions have continuous parameters, rank-clusters may be estimated by employing parameter fusion or shrinkage. Parameter fusion is the process of simultaneously estimating parameter values and groups of parameters that should be set equal in value (i.e., “fusing” parameters together). Masarotto and Varin (2012) analyze pairwise comparison data from sports tournaments with parameter fusion techniques under the Bradley-Terry model. Masarotto and Varin (2012) estimate an overall ranking of teams with rank-clusters by applying the frequentist fused lasso (Tibshirani et al., 2005), in which the absolute difference between every pair of worth parameters is penalized after some data-driven normalization. In this approach, the fused parameters are made equal and thus create a rank-cluster among the corresponding objects. The approach of Masarotto and Varin (2012) was extended to additional datasets in sports (Tutz and Schauberger, 2015) and academic journal rankings (Varin et al., 2016; Vana et al., 2016). Jeon and Choi (2018) argued that shrinkage methods like those proposed by Masarotto and Varin (2012) and Tutz and Schauberger (2015) were developed specifically for pairwise comparisons, and thus have inappropriate penalty functions for application to richer kinds of ordinal data like partial or complete rankings. As a result, Jeon and Choi (2018) proposed a modified regularization penalty that may be applied to partial or complete rankings under the Plackett-Luce model.
@@
Parameter fusion methods for rank-clustering exhibit five distinct disadvantages: First, maximum likelihood estimation of models in the BTL family, even in their simplest forms, often suffers from numerical instability and slow computational speed. As a result, numerous authors have proposed complex algorithms to improve estimation accuracy or speed (Hunter et al., 2004; Maystre and Grossglauser, 2015; Turner et al., 2020; Nguyen and Zhang, 2023). Second, uncertainty quantification is challenging and theoretically tenuous in lasso-based methods (Tibshirani, 1996; Fan and Li, 2001). Third, lasso penalty parameters may be difficult to select, requiring data-driven or ad hoc techniques (Tibshirani, 1996; Masarotto and Varin, 2012). Thus, interpretation of the resulting parameter estimates and associated uncertainty is reliant on the specific choice of penalty parameter. Fourth, prior knowledge on the amount and size of rank-clusters cannot be directly incorporated into the frequentist framework: Although the penalty parameter influences estimation of rank-clusters, the specific meaning of various possible choices is not directly interpretable. Fifth, to our knowledge only Jeon and Choi (2018) have formulated a parameter fusion-based method for ordinal data types beyond simple pairwise comparisons. As a result, only their parameter fusion method is available to estimate rank-clusters from complete or partial ranking data.
@@
Many of these disadvantages may be addressed using Bayesian methods. For example, Bayesian counterparts of the lasso and fused lasso—the Bayesian lasso (Park and Casella, 2008) and Bayesian fused lasso (Casella et al., 2010)—allow for valid uncertainty quantification and incorporation of prior knowledge via the selection of interpretable hyperparameters, as opposed to opaque penalty parameters. Bayesian methods for shrinkage and variable selection can be broadly categorized into two classes, continuous shrinkage and spike-and-slab priors. Examples of continuous shrinkage priors include the Bayesian Lasso (i.e., Laplace; Park and Casella (2008)), Bayesian Fused Lasso (Casella et al., 2010), t-distribution (Song and Cheng, 2020), Normal-Exponential-Gamma (Griffin and Brown, 2005; Shimamura et al., 2019), Normal-Gamma (Griffin and Brown, 2010), horseshoe (Carvalho et al., 2010), and Dirichlet-Laplace (Bhattacharya et al., 2015). No continuous shrinkage priors place positive probability on coefficients (or their differences) being precisely zero, which would encourage sparsity. Thus, parameter fusion must be performed via thresholding the posterior distribution, which is often ad-hoc (Porwal and Rodriguez, 2021). This leads us to consider priors of the latter class. Spike-and-slab priors (Mitchell and Beauchamp, 1988; George and McCulloch, 1997; Ishwaran and Rao, 2005) assign weight to both a point-mass at 0 (“spike”) and a continuous density function (“slab”). Although the specific formulations of these priors vary, they allow us to estimate which parameters are precisely zero in a probabilistic framework. Thus, spike-and-slab priors can be used for parameter fusion and uncertainty quantification. However, we are aware of only one existing modification of this prior class for parameter fusion: Wu et al. (2021) apply spike-and-slab to the differences in successive parameters in the linear regression setting. In their method, the order of parameters from least to greatest in coefficient value must be known in advance (as in the fused lasso). This is not practical in the ordinal data setting because the parameter order is equivalent to the overall ranking, whose estimation is a primary goal in the canonical preference data problem. Thus, no Bayesian parameter fusions methods exist which may be directly applied to ordinal data analyses with rank-clustering.
@@ 2.2 Methods based on Mallows Distributions
3 Methods
@@ 5 Applications
In this section, we apply our model to four real datasets involving ordinal comparisons. These four applications were chosen to highlight the applicability of our method to various ordinal data types and domain areas, which are summarized in Table 2. The data sets are comprised of sushi preferences of Japanese adults (Kamishima, 2003), policy preferences of respondents from Great Britain in a Eurobarometer survey (Reif and Melich, 1993), ranked-choice votes in a Minneapolis mayoral election (Minneapolis Elections and Voter Services, 2021), and pairwise game outcomes among teams in the National Basketball Association (National Basketball Association, 2021).
Section Setting Methodological Value
5.1 Sushi preferences in Tohoku Rank-clustering with complete rankings; comparing inferred overall ranking with that of the Clustered Mallows Model.
5.2 Minneapolis mayoral election votes Rank-clustering with top-3 partial rankings among 17 candidates; interpretable overall ranking captures the winner’s mandate in ranked-choice elections; comparing inferred overall ranking with those from a standard BTL model and two election procedures.
5.3 Eurobarometer survey policy preferences Rank-clustering with partial rankings; inferred overall ranking permits indistinguishability among options to aid policymakers.
5.4 Basketball game outcomes Rank-clustering with pairwise comparison data; inferred overall order of captures similarly-performing teams.
Table 2: Summary of applications.
@@ 5.1 Sushi Preferences in Tohoku
We first study complete preference rankings of 10 common sushi types based on a benchmarking dataset by Kamishima (2003). To allow our results to be comparable with an analysis of the sushi data in the related work of Piancastelli and Friel (2024), we analyze the preferences of survey respondents who lived in Japan’s Tohoku region until at least 15 years of age. In total, there were 280 respondents providing complete rankings of 10 sushi types. We fit a Rank-Clustered BTL distribution to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering but permit a wide variety of outcomes. Figure 5 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 5: Primary results of an analysis of preference rankings of 10 common sushi types in a survey of 280 Japanese adults. Posterior rank-clustering probabilities (left) and posterior distributions of sushi-specific worth parameters (right).
Sushi types are ordered according to their posterior median worth. Based on the left panel in Figure 5, fatty tuna appears to be the strictly most preferred sushi type based on this population, followed by tuna and shrimp rank-clustered with high posterior probability in second place. Salmon roe and sea eel exhibit high posterior probability of rank-clustering, as do sea urchin, tuna roll, and squid; these two groups may themselves be rank-clustered. Egg and cucumber roll are rank-clustered in last place. Our results demonstrate the proposed model’s ability to rank-cluster objects with uncertainty under complete rankings in survey data.
@@
Our second example analyzes real rank-choice votes from the 2021 mayoral election in Minneapolis, Minnesota (Minneapolis Elections and Voter Services, 2021). This election included 17 candidates (excluding write-ins and one who received no votes) and asked voters to rank their top-3 choices, in order. A total of 145,337 votes were cast in this election. To mimic exit polling data, we randomly sample 1000 valid votes for analysis, which we treat as a random sample of preferences from the population of Minneapolis voters. We want to estimate the overall preferences of Minneapolis voters regarding mayoral candidates and learn which candidates, if any, are rank-clustered at the population level. Clustering candidates may be of interest to political scientists or local political organizations for the purpose of understanding voter preferences (Gunther and Diamond, 2003; Dimock et al., 2014). For example, if the winner of the election is deemed to be rank-clustered with other candidate(s), their mandate may be considered weak. Conversely, if the winner is a singleton first-place rank-cluster—clearly ranked above all other candidates—their mandate may be considered strong. We fit a Rank-Clustered BTL to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage few of rank-clusters. Figure 6 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 6: Primary results of an analysis of 1000 valid votes in the 2021 Minneapolis mayoral election. Posterior rank-clustering probabilities (left) and posterior distributions of candidate-specific worth parameters (right).
In Figure 6, candidates are ordered by their posterior median estimate of worth. Cluster 1 consists of Jacob Frey, the winner and incumbent. We note that Frey is not rank-clustered with other candidates with high posterior probability, suggesting a relatively strong mandate. Cluster 2 consists of Kate Knuth and Sheila Nezhad, both non-incumbent DFL candidates with substantial political backing and fundraising. Clusters 3 and 4 consist of AJ Awed and Clint Conner, respectively, who are each DFL candidates with less experience and political backing than candidates in previous rank-clusters. Cluster 5 consists of Laverne Turner, a GOP candidate, and Cluster 6 consists of Carney Jr., Harcus, Globus, and Nelson. Last, Cluster 7 consists of 6 candidates with minimal support.
Figure 7 compares point estimates of rank for each candidate among four methods.
@@
The first and second rows display assigned ranks from ranked choice and “first-past-the-post” (FPP) election procedures, respectively. [^3] Since ranked choice and FPP elections use deterministic procedures, we use the term “assigned rank”. We calculate FPP ranks by ordering candidates by the number of first place votes he/she received (ignoring all second and third place votes). [^4] If the actual election had utilized FPP tabulation, results may have been different based on the differing voter strategies encouraged by ranked choice and FPP elections. The third and fourth rows display estimated ranks from a standard Bayesian BTL and our Rank-Clustered BTL, respectively. Estimated ranks are based on maximum a posteriori estimates. Frey wins the election in all methods. The BTL and Rank-Clustered BTL models roughly reflect the deterministic algorithms, although we notice some swaps in candidate ranks which may be attributed to differences between first place and second or third place votes. For example, Conner received fewer first place votes than Turner, but far more second and third place votes. As a result, deterministic algorithms rank Turner above Conner, while the BTL model takes into account the additional preference information and ranks Conner above Turner. In summary, the overall ordering estimated by the Rank-Clustered BTL differs from a standard BTL model and two deterministic election procedures. Furthermore, our model confirms that Frey is strictly preferred over the remaining candidates by voters.
@@
We subset the data to respondents from Great Britain to avoid heterogeneity and non-proportional sampling among respondents from different European countries. There were 1005 valid responses among this group (out of 1031 total surveyed), of which 970 were complete rankings and the rest ranked between 1 and 5 items (a top-6 ranking is inherently equivalent to a complete ranking since all survey options were presented). We seek to identify a population-level ordering of the priorities that accounts for potential equality or indistinguishability among the options based on the survey data. We note that these data were previously studied by Wang et al. (2017) in a mixed-membership model. Our analysis, although a simplification of the diverse population’s heterogeneous preferences, provides a simpler interpretation to policy-makers interested in understanding holistic attitudes of their constituent population.
@@
Figure 8: Primary results of an analysis of respondents from Great Britain to question 28 in the Eurobarometer 34.1 survey. Posterior rank-clustering probabilities (left) and posterior distributions of survey option-specific parameters (right).

### v2

@@ Abstract
Traditional statistical inference on ordinal comparison data results in an overall ranking of objects, e.g., from best to worst, with each object having a unique rank. However, ranks of some objects may not be statistically distinguishable. This could happen due to insufficient data or to the true underlying object qualities being equal. Because uncertainty communication in estimates of overall rankings is notoriously difficult, we take a different approach and allow groups of objects to have equal ranks or be rank-clustered in our model. Existing models related to rank-clustering are limited by their inability to handle a variety of ordinal data types, to quantify uncertainty, or by the need to pre-specify the number and size of potential rank-clusters. We solve these limitations through our proposed Bayesian Rank-Clustered Bradley-Terry-Luce model. We accommodate rank-clustering via parameter fusion by imposing a novel spike-and-slab prior on object-specific worth parameters in Bradley-Terry-Luce family of distributions for ordinal comparisons. We demonstrate rank-clustering on simulated and real datasets in surveys, elections, and sports analytics.
@@ 1 Introduction
In a traditional analysis of ordinal data, we assume $I$ judges assess $J$ objects by providing ordinal preferences, $\Pi$. Each judge’s ordinal preferences, $\Pi_{i}$, may be provided in various forms, such as complete rankings, partial rankings, or pairwise comparisons among the available objects or some subset thereof. Standard statistical model families for ranking data such as Mallows (Mallows, 1957) or Bradley-Terry-Luce (Bradley and Terry, 1952; Plackett, 1975; Luce, 1959) derive or estimate the rank of each object whereby each object receives a unique rank. An estimated overall ranking then orders all objects from best to worst. Analyses of this kind are used to rank candidates in ranked choice elections, (Gormley and Murphy, 2008; Mollica and Tardella, 2017), sports teams or players in a league using pairwise game outcomes (Tutz and Schauberger, 2015; Barrientos et al., 2023), or genes based on ordinal comparisons of genomics data (Eliseussen et al., 2023; Vitelli et al., 2018).
@@
However, requiring estimated ranks to be unique is not always useful or appropriate. For example, some objects may be equal or indistinguishable in their true quality or ability. Consider an election in which 2 candidates, both of the same political party, are running for an office. If voters express their preferences solely on the basis of party, the candidates are inherently equal in quality. In another situation, when the number of votes cast is small, estimated ranks assigned to each candidate could exhibit substantial uncertainty, suggesting the candidates are indistinguishable in quality based on the limited number of observed votes. In such situations, allowing for inference to estimate the candidates as having the same rank or be rank-clustered may improve interpretability, prediction, and decision-making when analyzing ordinal preferences.
@@
In this paper, we propose a Bayesian framework for ordinal data analysis that estimates an overall ranking of objects with rank-clusters, develop a computationally-efficient Gibbs sampler for estimation, and apply the model to real and simulated data. Specifically, we choose to model observed rank via the Bradley-Terry-Luce (BTL) family of distributions which permits analysis of ordinal preferences in many forms, such as complete rankings, partial rankings, pairwise comparisons, and groupwise comparisons. To induce rank-clusters, we place a novel spike-and-slab fusion prior on the object-specific parameters of BTL distributions. In contrast to existing work related to rank-clustering in the literature, our model requires neither the parameter order nor the number or size of rank-clusters to be known in advance. Instead, these quantities are treated as random variables and estimated simultaneously so that their corresponding uncertainty is naturally reflected in the resulting inferences.
@@
The rest of the paper is organized as follows. We first review literature related to rank-clustering in Section 2. Then, we propose the Partition-based Spike-and-Slab Fusion prior and apply it to a BTL model for ordinal data in Section 3. We develop a computationally-efficient Gibbs sampler based on reversible jump Markov chain Monte Carlo and demonstrate its accuracy on simulated data in Section 4. To demonstrate a wide variety of methodological benefits of our proposed framework, in Section 5, we apply the model to four real datasets: (i) complete rankings of sushi preferences provided by Japanese adults, (ii) partial rankings of 2021 Minneapolis mayoral candidates expressed by voters in a ranked choice election, (iii) complete and partial rankings of policy options from Eurobarometer 34.1, a survey which measures various European attitudes, and (iv) pairwise basketball game outcomes from the 2023-24 season of the National Basketball Association. We conclude with a brief discussion in Section 6.
@@ 2 Background
Before reviewing ordinal comparisons literature, it is helpful to introduce some basic terminology and notation. Rankings are a type of ordinal preferences that denote a relative ordering of objects from best to worst, potentially allowing ties. We use the operator $\prec$ to denote a strict ordering of two objects; e.g., $A\prec B$ states that object $A$ is strictly preferred to $B$. An object’s rank is the place it receives in the ranking. [^1] Although some authors have drawn a distinction between the terms “ranking” and “ordering,” in this paper we choose to use solely the former in accordance with its popular usage. Rankings arise in different forms. Given a collection of objects, a ranking is called complete when all objects are ranked. In contrast, a ranking is called partial when all objects were considered, but only a subset of the most-preferred are ranked (e.g., a top-$5$ ranking). In a partial ranking, we assume that unranked objects are less-preferred than those ranked, but also that the preference order among the unranked objects is unknown. Next, we call a ranking incomplete when a judge is asked only to rank a subset of the complete collection of objects. In incomplete rankings, no information can be gleaned regarding objects not considered. For example, if a voter is asked by an election pollster to rank candidates from a single political party, the ranking should provide no information regarding their preferences on candidates from other parties. We call incomplete rankings involving two objects (candidates in the above example) a pairwise comparison, and incomplete rankings involving more than two objects a groupwise comparison. Rankings may be both partial and incomplete; e.g., it could be a top-3 ranking of mayoral candidates from a specific political party.
Next, we briefly review methods for estimating rank-clusters based on the Bradley-Terry-Luce and Mallows families of ordinal data models in turn. For a more thorough review of these standard model families, see Marden (1996) and Alvo and Yu (2014).
@@ 2.1 Methods based on Bradley-Terry-Luce Distributions
Since BTL distributions have continuous parameters, rank-clusters may be estimated by employing parameter fusion or shrinkage. Parameter fusion is the process of simultaneously estimating parameter values and groups of parameters that should be set equal in value (i.e., “fusing” parameters together). Masarotto and Varin (2012) analyze pairwise comparison data from sports tournaments with parameter fusion techniques under the Bradley-Terry model. Masarotto and Varin (2012) estimate an overall ranking of teams with rank-clusters by applying the frequentist fused lasso (Tibshirani et al., 2005), in which the absolute difference between every pair of worth parameters is penalized after some data-driven normalization. In this approach, the fused parameters are made equal and thus create a rank-cluster among the corresponding objects. The approach of Masarotto and Varin (2012) was extended to additional datasets in sports (Tutz and Schauberger, 2015) and academic journal rankings (Varin et al., 2016; Vana et al., 2016). Jeon and Choi (2018) argued that shrinkage methods like those proposed by Masarotto and Varin (2012) and Tutz and Schauberger (2015) were developed specifically for pairwise comparisons, and thus have inappropriate penalty functions for application to richer kinds of ordinal data like partial or complete rankings. As a result, Jeon and Choi (2018) proposed a modified regularization penalty that may be applied to partial or complete rankings under the Plackett-Luce model. Relatedly, Hermes et al. (2024) consider sparse estimation of a Plackett-Luce model with object-level covariates under judge heterogeneity. In their setting, the number of heterogeneous preference groups and the group membership of each judge are assumed fixed and known. To improve efficiency of estimation across groups and predictive performance, they impose a lasso penalty on group-specific covariate coefficients and a simultaneous fused lasso penalty between each pair of group-specific covariate coefficients. We note that the setting studied by Hermes et al. (2024) is fundamentally different to ours, in that they assume (known) preference heterogeneity among the judges and the presence of object-specific covariates.
@@
Parameter fusion methods for rank-clustering exhibit four distinct disadvantages: First, maximum likelihood estimation of models in the BTL family, even in their simplest forms, often suffers from numerical instability and slow computational speed. As a result, numerous authors have proposed complex algorithms to improve estimation accuracy or speed (Hunter et al., 2004; Maystre and Grossglauser, 2015; Turner et al., 2020; Nguyen and Zhang, 2023). Second, uncertainty quantification is challenging and theoretically tenuous in lasso-based methods (Tibshirani, 1996; Fan and Li, 2001). Third, lasso penalty parameters may be difficult to select, requiring data-driven or ad hoc techniques (Tibshirani, 1996; Masarotto and Varin, 2012). Thus, interpretation of the resulting parameter estimates and associated uncertainty is reliant on the specific choice of penalty parameter. Fourth, prior knowledge on the amount and size of rank-clusters cannot be directly incorporated into the frequentist framework: Although the penalty parameter influences estimation of rank-clusters, the specific meaning of various possible choices is not directly interpretable.
@@
Many of these disadvantages may be addressed using spike-and-slab priors, a Bayesian approach to variable selection (Mitchell and Beauchamp, 1988; George and McCulloch, 1997; Ishwaran and Rao, 2005). Spike-and-slab priors assign weight to both a point-mass at 0 (“spike”) and a continuous density function (“slab”). Although the specific formulations of these priors vary, they estimate parameters which are precisely zero in a probabilistic framework that incorporates prior knowledge via interpretable hyperparameters, as opposed to opaque penalty parameters. However, we are aware of only one variant of this prior class for parameter fusion: Wu et al. (2021) apply spike-and-slab to differences in successive parameters in a linear regression. In their method, the order of parameters from least to greatest in coefficient value must be known in advance (as in the fused lasso). This is not practical in the canonical ordinal data setting because the parameter order is equivalent to the overall ranking, whose estimation is a primary goal. Thus, no Bayesian parameter fusions methods exist which may be directly applied to ordinal data analyses with rank-clustering. Alternatively, one may consider the class of continuous shrinkage priors, which include Bayesian variants of the lasso (Park and Casella, 2008) and fused lasso (Casella et al., 2010) among others (e.g., Griffin and Brown (2005); Carvalho et al. (2010); Bhattacharya et al. (2015)). However, continuous shrinkage priors do not place positive probability on coefficients (or their differences) being precisely zero. Thus, parameter fusion must be performed via thresholding the posterior distribution, which is often ad-hoc (Porwal and Rodriguez, 2021) and will not be considered in this work.
@@ 2.2 Methods based on Mallows Distributions
3 The Rank-Clustered Bradley-Terry-Luce Model
@@ 5 Applications
In this section, we apply the Rank-Clustered BTL model to four real datasets involving ordinal comparisons. These four applications were chosen to highlight the applicability of our method to various ordinal data types and domain areas and illustrate methodological values of our approach which are summarized in Table 2. The data sets are comprised of sushi preferences of Japanese adults (Kamishima, 2003), policy preferences of respondents from Great Britain in a Eurobarometer survey (Reif and Melich, 1993), ranked-choice votes in a Minneapolis mayoral election (Minneapolis Elections and Voter Services, 2021), and pairwise game outcomes among teams in the United States National Basketball Association (National Basketball Association, 2024).
Setting Data Type Methodological Value
5.1 Sushi preferences in Tohoku Complete rankings of 10 sushi types Rank-clusters sushi types by preferences. Comparing inferred overall ranking with that of the Clustered Mallows Model.
5.2 Minneapolis mayoral election votes Top-3 partial rankings of 17 candidates Interpretable overall ranking captures the winner’s mandate in ranked-choice elections. Comparing inferred overall ranking with those from a standard BTL model and two election procedures.
5.3 Eurobarometer survey policy preferences Partial rankings of 7 policy options Inferred overall ranking permits identification of similarly preferred options to aid policymakers.
5.4 Basketball game outcomes Pairwise comparisons (game winners) among 30 teams Inferred overall order captures similarly-performing teams. Setting with limited information and low signal.
Table 2: Summary of applications by subsection.
@@ 5.1 Sushi Preferences in Tohoku
We first study complete preference rankings of 10 sushi types from a benchmarking dataset by Kamishima (2003). To allow our results to be comparable with an analysis of the sushi data by Piancastelli and Friel (2024), we analyze the preferences of survey respondents who lived in Japan’s Tohoku region until at least 15 years of age. There were 280 such respondents. We fit a Rank-Clustered BTL distribution to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering but permit a wide variety of outcomes. Figure 5 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics are provided in the supplementary materials.
Figure 5: Primary results from Rank-Clustered BTL analysis of Tohoku sushi data. Left: Posterior rank-clustering probabilities. Main diagonal displays posterior median estimate of worth parameter after normalization. Red squares indicate maximum a posteriori rank-clusters. Right: Posterior distributions of sushi-specific worth parameters.
Sushi types are ordered according to posterior median worth. Based on the left panel in Figure 5, fatty tuna appears to be strictly most preferred in this population, followed by tuna and shrimp rank-clustered in second place. Salmon roe and sea eel exhibit high posterior probability of rank-clustering, as do sea urchin, tuna roll, and squid; these two groups may themselves be rank-clustered. Egg and cucumber roll are rank-clustered in last place. Our results demonstrate the proposed model’s ability to rank-cluster objects with uncertainty under complete rankings in survey data.
@@
Our second example analyzes real rank-choice votes from the 2021 mayoral election in Minneapolis, Minnesota (Minneapolis Elections and Voter Services, 2021). This election included 17 candidates (excluding write-ins and one who received no votes) and asked voters to rank their top-3 choices, in order. A total of 145,337 votes were cast in this election. To mimic exit polling data, we randomly sample 1000 valid votes for analysis, which we treat as a random sample of preferences from the population of Minneapolis voters. We want to estimate the overall preferences of Minneapolis voters regarding mayoral candidates and learn which candidates, if any, are rank-clustered at the population level. Clustering candidates may be of interest to political scientists or local political organizations for the purpose of understanding voter preferences (Gunther and Diamond, 2003; Dimock et al., 2014). For example, if the winner of the election is deemed to be rank-clustered with other candidate(s), their mandate may be considered weak. Conversely, if the winner is a singleton first-place rank-cluster—clearly ranked above all other candidates—their mandate may be considered strong. We fit a Rank-Clustered BTL to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage few rank-clusters. Figure 6 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics are provided in the supplementary materials.
Figure 6: Primary results from Rank-Clustered BTL analysis of mayoral votes. Left: Posterior rank-clustering probabilities. Main diagonal displays posterior median estimate of worth parameter after normalization. Red squares indicate maximum a posteriori rank-clusters. Right: Posterior distributions of candidate-specific worth parameters.
In Figure 6, candidates are ordered by their posterior median estimate of worth. Cluster 1 consists of Jacob Frey, the winner and incumbent. We note that Frey is not rank-clustered with other candidates with high posterior probability, suggesting a relatively strong mandate. Cluster 2 consists of Kate Knuth and Sheila Nezhad, both female, non-incumbent DFL candidates. Last, Cluster 7 consists of 6 candidates with minimal support.
Figure 7 compares point estimates of rank for each candidate across four methods.
@@
The first and second rows display assigned ranks from ranked choice and “first-past-the-post” (FPP) election procedures, respectively. We calculate FPP ranks by ordering candidates by the number of first place votes he/she received (ignoring all second and third place votes). [^3] If the actual election had utilized FPP tabulation, results may have been different based on the differing voter strategies encouraged by ranked choice and FPP elections. The third and fourth rows display maximum a posteriori ranks from a standard Bayesian BTL and our Rank-Clustered BTL, respectively. Frey wins the election in all methods. The BTL and Rank-Clustered BTL models roughly reflect the deterministic algorithms, although we notice some swaps in candidate ranks which may be attributed to differences between first place and second or third place votes. For example, Conner received fewer first place votes than Turner, but far more second and third place votes (see supplementary materials for vote totals). As a result, deterministic algorithms rank Turner above Conner, while the BTL model takes into account the additional preference information and ranks Conner above Turner. In summary, the overall ordering estimated by the Rank-Clustered BTL differs from a standard BTL model and two deterministic election procedures. Furthermore, our model confirms that Frey is strictly preferred over the remaining candidates by voters.
@@
We subset the data to respondents from Great Britain to avoid heterogeneity and non-proportional sampling among respondents from different European countries. There were 1005 valid responses among this group (out of 1031 total surveyed), of which 970 were complete rankings and the rest ranked between 1 and 5 items (a top-6 ranking is inherently equivalent to a complete ranking since all survey options were presented). We seek to identify a population-level ordering of the priorities that accounts for potential equality or indistinguishability among the options based on the survey data. These data were previously studied by Wang et al. (2017) with a mixed-membership model to learn about heterogeneity of opinions among survey respondents. Our analysis, although a simplification of the diverse population’s heterogeneous preferences, provides a simpler interpretation to policy-makers interested in understanding rank-ordering of policy preferences.
@@
Figure 8: Primary results from Rank-Clustered BTL analysis of Eurobarometer 34.1 data. Left: Posterior rank-clustering probabilities. Main diagonal displays posterior median estimate of worth parameter after normalization. Red squares indicate maximum a posteriori rank-clusters. Right: Posterior distributions of policy-specific worth parameters.